# Computation of $\left\langle\Phi^{2}\right\rangle$ and quantum fluxes at the polar interior of a spinning black hole

### v1

@@ D.3.2 Regularization of the $\omega$ integral
In addition to this $i\geq 1$ spectrum of oscillations, we also find another (single) oscillation with a different exponential factor and a significantly longer $\omega$-wavelength, hence we denote it by $i=0$. The origin of this longer oscillation is not entirely clear to us. Nevertheless, a simple analysis, based on a somewhat speculative argument suggests that the parameters characterizing the oscillation are given by
@@
$\Phi_{0}\left(r\right)=\left[\frac{1}{2\kappa_{+}}\ln\left(a^{2}+r_{+}^{2}\right)-\frac{1}{2\kappa_{-}}\ln\left(a^{2}+r_{-}^{2}\right)-4M\ln\left(r_{+}-r_{-}\right)\right]-2r_{*}\left(r\right)\,.\$ (166)
@@
The oscillation cancellation procedure, which accommodates an exponentially growing amplitude, constitutes of taking averages with suitable weights. In our case, the suitable self cancellation operator for a function $f\left(x\right)$ that behaves like $\propto e^{\eta x}e^{(2\pi x)i/\tilde{\lambda}}$ is
@@ D.5 Numerical results and comparison with the standard variant
Figure 19: The analytic-extension integrand $T_{uu}^{\text{ae}}(\omega)$ at $r=1.5M$ within a BH of spin $a/M=0.8$, before and after its regularization including the required oscillation cancellation (as described in the main text). Left: the basic integrand on a logarithmic scale, where the dominant exponent with the growth parameter given in Eq. (164) is clearly visible alongside oscillations. Right: the integrand exposed after regularization, to be integrated in the subsequent step. While the details of the applied operations affect the appearance of the resulting integrand shown in this panel, the final integral value remains invariant (refer to Footnote 44). The comparison between the two panels demonstrates the success of the damping procedure, which is vital in order to reveal the ‘real’ physical content of the modes.

### v2

@@ D.3.2 Regularization of the $\omega$ integral
In addition to this $i\geq 1$ spectrum of oscillations, we also find another (single) oscillation with a different exponential factor and a significantly longer $\omega$-wavelength, hence we denote it by $i=0$. The origin of this longer oscillation is not entirely clear to us. Nevertheless, a simple analysis, based on a somewhat speculative argument (related to a special type of connecting null geodesics  [^45] Specifically, this is a polar (i.e. $\theta=0$) complex null geodesic that connects the $r$ value of the evaluation point with the complex point $r=ia$ — which is the location of the pole in the effective potential $V_{\omega lm}(r)$ give in Eq. (22). The (complex) t-difference $\Delta t$ between these two edge points of the geodesic can be easily computed analytically; and when translated to an $\omega$-frequency (and hence to corresponding $\eta$ and $\Phi$ parameters according to the connection that was briefly described above), one obtains the values given in Eqs. (165,166).) suggests that the parameters characterizing the oscillation are given by  [^46] Note, however, that in order to correctly perform the oscillation cancellation, analytical determination of the oscillation parameters $\eta$ and $\Phi$ is not necessary (these parameters can be determined numerically).:
@@
$\Phi_{0}\left(r\right)=\left[\frac{1}{2\kappa_{+}}\ln\left(a^{2}+r_{+}^{2}\right)-\frac{1}{2\kappa_{-}}\ln\left(a^{2}+r_{-}^{2}\right)-4M\ln\left(r_{+}-r_{-}\right)\right]-2r_{*}\left(r\right)\,,\$ (166)
@@
The oscillation cancellation procedure, which accommodates an exponentially growing amplitude, constitutes of taking averages with suitable weights (which sum to 1). In our case, the suitable self cancellation operator for a function $f\left(x\right)$ that behaves like $\propto e^{\eta x}e^{(2\pi x)i/\tilde{\lambda}}$ is
@@ D.5 Numerical results and comparison with the standard variant
Figure 19: The analytic-extension integrand $T_{uu}^{\text{ae}}(\omega)$ at $r=1.5M$ within a BH of spin $a/M=0.8$, before and after its regularization including the required oscillation cancellation (as described in the main text). Left: the basic integrand on a logarithmic scale, where the dominant exponent with the growth parameter given in Eq. (164) is clearly visible alongside oscillations. Right: the integrand exposed after regularization, to be integrated in the subsequent step. While the details of the applied operations affect the appearance of the resulting integrand shown in this panel, the final integral value remains invariant (refer to footnote 47). The comparison between the two panels demonstrates the success of the damping procedure, which is vital in order to reveal the ‘real’ physical content of the modes.